OpenStudy (anonymous):

How many milligrams of NaOH are required to prepare 10.50dL of solution?

OpenStudy (anonymous):

Am I on the right track so far? N=14.01g O=16.00g H=1.008g NaOH=31.02g

OpenStudy (anonymous):

\[10.50dL \times \frac{ 0.10L }{ 1dL} \times \frac{ 1000g }{ 1L } \times \frac{ 1000mg }{ 1g}\] Milligrams of NaOH before solution=Milligrams of in Solution 0.0

OpenStudy (anonymous):

Milligrams of NaOH*

OpenStudy (anonymous):

Thanks so much, I completely blanked!

OpenStudy (anonymous):

No problem! Good luck on the rest.